what will be volume if we convert 20 percent 20 ml sugar solution into 15 percent sugar solution ??
Answer by lwsshak3(11628)

what will be volume if we convert 20 percent 20 ml sugar solution into 15 percent sugar solution ??
**
let x=increase in volume
20%*20=15%(20+x)
4=3+.15x
.15x=1
x=1/.15= 6-2/3
new volume=20+x=26-2/3 ml